if 6 women & 10 children completes the work in 6 days , and 6 Women alone completes the same work in 10 days then how many days does 10 children will take to complete the same work?
Answer by ptaylor(2198)

Let x=number of days it takes 10 children to complete the work
Then 10 children work at the rate of 1/x of the work per day
6 women and 10 children work at the rate of 1/6 of the work per day
6 women alone works at the rate of 1/10 of the work per day
So, our equation to solve is:
1/10 + 1/x = 1/6 multiply each term by 30x
3x+30=5x
2x=30
x=15 days ---time it takes 10 children working alone to complete the work
CK
1/10 + 1/15=1/6
3/30 +2/30=1/6
5/30=1/6
1/6=1/6
Another way to check:
In 6 days, 6 women do 6/10 of the work
In 6 days 10 children do 6/15 of the work
6/10 + 6/15 should add up to 1 (all the work)
18/30 + 12/30=30/30

Hope this helps---ptaylor